Circular No. 6744

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

IRAS 18119–1342 A. Henden, Universities Space Research Corporation (USRA) and U.S.

Naval Observatory, Flagstaff Station (USNOFS); C. Luginbuhl and F. Vrba, USNOFS; B. Canzian, USRA and USNOFS; J. Munn, USNOFS; and D. Hartmann, Clemson University, write: “We report a strong near-infrared source at α = 18^{h}14^{m}50^{s}.31, δ = −13^{o}41′05′′.6 (equinox 2000.0, ± 0′′.5), which is ∼ 2′ from the center of the IPN (cf. IAUC 6593, 6683) error box for the soft γ-ray repeater reported on IAUC 6743. JHK images saturate in our shortest available exposure (0.16 s), but we estimate magnitudes J ∼ 7.4, H ∼ 6.0, K ∼ 5.1. This object is coincident with IRAS 18119–1342, within the errors of the positions. CCD images with limiting magnitudes around 20–21 show the object barely visible at V but easily visible at R and bright at I. The object is barely on the POSS-I red plate. While the GRB error-box center has been recently revised, based on RXTE ASM data (D. Smith, private comm.), to a point some 20′ south of the IPN center, we note the similarity of this source to the potential infrared counterparts of SGR 1900+14.”

GRB 970815 A. J. Castro-Tirado and J. Gorosabel, Laboratorio de Astrofísica Es-

pacial y Física Fundamental (LAEFF), Madrid; and J. Greiner, Astrophysikalisches Institut, Potsdam, communicate: “Following the detection of transient x-ray emission from GRB 970815 by RXTE (IAUC 6718), frames in the K′ band were obtained by A. Campos (Observatorio Astronómico Nacional, Alcalá de Henares) and N. Metcalfe (Durham University) on Aug. 17 at the 3.5-m telescope (+ OMEGA) at the German-Spanish Calar Alto Observatory (CAHA). R and B-band images were acquired on Aug. 20 by E. Sanchez (Instituto de Astrofísica de Andalucía, Granada) with the 1.5-m telescope at the Sierra Nevada Observatory. Additional images were obtained by H.-J. Hagen (Hamburg Observatory) with the 2.2-m CAHA telescope (+ CAFOS) on Aug. 20–21 and by J. Trapero (LAEFF) with the 4.2-m William Herschel Telescope at La Palma on Aug. 22. No source is detected within the 10′′ radius error box of the faint ROSAT x-ray source on the border of the RXTE region (IAUC 6742). Upper limits are: Aug. 17, K′ > 18; Aug. 20, B > 21.5, R > 22; Aug. 23, B > 22.5, R > 23. Deeper optical observations are encouraged.”

1997 September 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green